OpenStudy (eujc21):

Find the length of the following curve: r = 1, /theta = pi/6 , 0 < phi < pi/3?

OpenStudy (eujc21):

\[r = 1, \theta = \frac{ \pi }{ 6 } , 0 < \phi < \frac{ \pi }{ 3 }\]

OpenStudy (anonymous):

Is this all the question says? There is only one point with both \(r=1, \theta = \frac{\pi}{6}\), or are those supposed to be two separate lines? And the \(\phi\) bit makes no sense either. (What is \(\phi\)?)

OpenStudy (dumbcow):

yeah i was going to help earlier but could you clarify the problem? what does curve look like?

OpenStudy (eujc21):

Well this is all the problem states Find the length of the following curves. The first one was a) r=3, pi/4 < phi < pi/2, z = constant. (Cylindrical) So I used the arc length and neglected z. b) which is the current problem just states the beginning with the parameters i posted earlier for (Spherical)

OpenStudy (dumbcow):

oh ok thats kinda what i thought